An oil film of thickness 1.5 mm is used for lubrication in between a square plate of side 0.9 m into 0.9 m and and inclined plane having an angle of inclination 20
1
Expert's answer
2021-07-21T09:49:02-0400

We need to find the dynamic viscosity of the oil.

Now shear stress "τ = μ \frac{du }{ dy}"

Force / unit area = "μ \frac{du }{ dy}"

Now "F = 392.4 , A = 0.9 * 0.9 = 0.81 sq m, dy = 10 mm = 10 * 10^{-3} m"

"So \\
\frac{F}{A }= \frac{392.4 }{ 0.81} = 484.4\\
 So \space we \space get\\
 \frac{F}{A } = μ \frac{du}{ dy}\\
 484.4 = μ (\frac{0.2 }{ 0.01})\\
 484.4 = μ *20\\
 μ = \frac{484.4 }{ 20}\\ 
 μ = 24.22 Ns / m^2\\
 μ = 242.2 poise"